# On the virialization threshold for halo mass functions

### v1

@@ Appendix A $\delta_{{\rm v}}$ fit for $\Lambda$CDM and a general code for its computation
$\delta_{{\rm v}}=\alpha\delta_{{\rm c}}^{\beta}$ (A.1)
@@
$\beta=\frac{\ln\left(\delta_{\rm v}/\delta_{{\rm v}\text{EdS}}\right)}{\ln\left(\delta_{c}/\delta_{c\text{EdS}}\right)}\,.$ (A.2)
For $\Omega_{m}^{0}=0.3$ and $20<z<$ 2, we have $-2.27>\beta>-2.55$. We compute an effective value in the same way as in equation (3.11), getting
@@
In figure 8, we show the percent difference between the numerical solution and the fit given by (A.1). As can be seen, this representation of $\delta_{{\rm v}}$ is very precise and robust when varying $\Omega_{m}^{0}$, with error below $0.17\%$ for $0.40.2\leq\Omega_{m}^{0}<$ 0.4.
Figure 8: Percent difference between the representation of $\delta_{{\rm v}}$ given by (A.1) and its numerical solution in the $\Lambda$CDM model for several values of $\Omega_{m}^{0}$ indicated in the plot legend.

### v2

@@ Appendix A $\delta_{{\rm v}}$ fit for $\Lambda$CDM and a general code for its computation
$\delta_{{\rm v}}(z)=\alpha\delta_{{\rm c}}^{\beta}(z)$ (A.1)
@@
$\beta=\frac{\ln\left(\delta_{\rm v}(z)/\delta_{{\rm v}\text{EdS}}\right)}{\ln\left(\delta_{c}(z)/\delta_{c\text{EdS}}\right)}\,.$ (A.2)
For $\Omega_{m0}=0.3$ and $20<z<$ 2, we have $-2.27>\beta>-2.55$. We compute an effective value in the same way as in equation (3.11), getting
@@
In figure 7, we show the percent difference between the numerical solution and the fit given by (A.1). As can be seen, this representation of $\delta_{{\rm v}}$ is very precise and robust when varying $\Omega_{m0}$, with error below $0.17\%$ for $0.40.2\leq\Omega_{m0}<$ 0.4.
Figure 7: Percent difference between the representation of $\delta_{{\rm v}}$ given by (A.1) and its numerical solution in the $\Lambda$CDM model for several values of $\Omega_{m0}$ indicated in the plot legend.